# Partial Semigroupoid Actions on Sets

### v1

@@ Abstract.
We introduce partial semigroupoid actions on sets and demonstrate that each such action admits universal globalization. Our construction extends the universal globalization for partial category actions given by P. Nystedt (Lundström) and the tensor product globalization for strong partial semigroup actions given by G. Kudryavtseva and V. Laan, thereby unifying the theory of partial actions for both categories and semigroups.
@@ 1. Introduction
Partial group actions on sets were introduced by R. Exel in [11], where he characterizes a class of $C^{\ast}$-algebras as crossed products by partial actions. Examples of partial group action on sets can be easily obtained from a restriction of a (global) group action to a subset of a given set. In [1, Teorema 4.23], F. Abadie proved that any partial group action on a set can indeed be viewed as a restriction of a (unique) global one. This global action was later called universal globalization, due to its universal property. In [16, Theorem 3.4], J. Kellendonk and M. Lawson proved that the universal globalizations are initial objects in a determined category. Partial actions and the globalization problem were investigated in many other cases, such as partial actions of monoids ([15] and [20]), semigroups [17], ordered groupoids ([14] and [19]), groupoids [3], inverse categories [2], categories [21], and inverse semigroupoids [8]. Another contexts of partial actions also can be found in the surveys [5], [9] and [10], and in Exel’s book [13].
@@
As an algebraic structure, a semigroupoid is a set equipped with a partially defined associative operation. It is worth mentioning that there is another (more restrict) notion of semigroupoid due to B. Tilson [22]. The latter can be seen as a directed graph equipped with a partially defined associative operation and, therefore, it covers the class of categories and semigroups. However, the class of Exel’s semigroupoids is strictly larger than the class of Tilson’s semigroupoids (also referred to as semicategories or categorical semigroupoids). Several properties of semigroupoids and graphed semigroupoids can be found in the work written by L. Cordeiro in [6].
In this paper, we are interested in studying the globalization problem for partial semigroupoid actions on sets. Our method generalizes the work of P. Nystedt (Lundström) in [21], whose construction of a universal globalization for partial category actions traces back to F. Abadie’s approach in [1]. We also compare our results with the tensor product globalization for partial semigroup actions presented by G. Kudryavtseva and V. Laan in [17], concluding that our universal globalization coincides with the tensor product globalization when the partial semigroup action is non-degenerate.
The paper is structured as follows: in the second section we present the definition of semigroupoid and its partial actions, along with the construction of the standard restriction of a global action to a subset, and the definition of universal globalization. In the third section, we present the construction of a universal globalization for any given partial action. Finally, in the last section, we explore the relationship between our construction and those presented for categories and semigroups.
@@ 2. Global and Partial Actions
For our purposes, by a semigroupoid we mean an Exel’s semigroupoid, which is a more general definition compared to Tilson’s semigroupoid.
@@
In other words, $S^{s}$ is the set of elements in $S$ that are right composable with $s$, while ${}^{s}S$ is the set of elements in $S$ that are left composable with $s$. Due to the associativity of $S$, we have $S^{t}=S^{st}$ and ${{}^{s}S}={{}^{st}S}$, whenever $(s,t)\in S^{(2)}$. A semigroupoid is called categorical if, for every $s,t\in S$, the sets $S^{s}$ and $S^{t}$ are either disjoint or equal; equivalently, if ${{}^{s}S}$ and ${{}^{t}S}$ are either disjoint or equal.
The class of semigroupoids generalizes two relevant algebraic structures: the semigroups, which can be seen as the semigroupoids whose composition is totally defined; and categories, which can be seen as semigroupoids with local identities. We will present the definition of those classes for future reference.
@@
By definition, the semigroups are precisely the semigroupoids such that $S^{(2)}=S\times S$. Furthermore, every semigroup is a categorical semigroupoid, since $\mathcal{S}^{s}=\mathcal{S}^{t}$ for every pair $s,t\in\mathcal{S}$.
@@
Notice that, if $1_{e}\circ 1_{f}$ is defined, then $1_{e}=1_{f}$ and, therefore, $e=D(1_{e})=D(1_{f})=f$. Consequently, $e\mapsto 1_{e}$ is an injection and, hence, we can assume $\mathcal{C}_{0}\subseteq\mathcal{C}_{1}=:\mathcal{C}$. Furthermore, for every $f\in\mathcal{C}_{1}$, the composition $1_{R(f)}\circ f$ is defined. Therefore, if $f$ is an identity, it must be $f=1_{R(f)}\in\mathcal{C}_{0}$. Therefore, $\mathcal{C}_{0}$ is precisely the set of identities of $\mathcal{C}_{1}$.
A category can be seen as a semigroupoid such that $\mathcal{C}^{(2)}=\{(f,g)\colon D(f)=R(g)\}\subseteq\mathcal{C}\times\mathcal{C}$, and such that every element $f\in\mathcal{C}$ has a left identity $R(f)$ and a right identity $D(f)$. Furthermore, every category is a categorical semigroupoid. In fact, for every $f\in\mathcal{C}$ we have:
@@
[22, Appendix B] A graphed semigroupoid (also called Tilson semigroupoid) is a quintuple $(G_{0},G_{1},d,r,\star)$, where:
@@
Markov semigroupoids constitute a class of semigroupoids that contains the path categories of directed graphs and some semigroupoids that cannot be realized as graphed semigroupoids. They are constructed in the following way: Let $\Sigma$ be a set and $A=(A(x,y))_{x,y\in\Sigma}$ be a 0-1 matrix. Assume that:
@@
Let $\Sigma=\{s,t\}$ and $A$ be the $2\times 2$ matrix such that $A(s,t)=1$ and $A(x,y)=0$, for all $(x,y)\neq(s,t)$. The Markov semigroupoid associated to $(\Sigma,A)$ is the set $S=\{s,t,st\}$, where the only composition defined is $s\star t=st$. This Markov semigroupoid can be endowed with a graph structure $(S_{0},S_{1},d,r)$, in the following way: $S_{0}$ consists of three vertex, say $S_{0}=\{e_{1},e_{2},e_{3}\}$, $S_{1}=S$, and the functions $d,r\colon S\to S_{0}$ are given by $d(t)=d(st)=e_{1}$, $d(s)=r(t)=e_{2}$ and $r(s)=r(st)=e_{3}$.
@@
Furthermore, it is not a Markov semigroupoid. Specifically, we have $s_{1}\star t_{1}=s_{1}\star t_{2}$, while $t_{1}\neq t_{2}$ (note that any element in a Markov semigroupoid has a unique factorization into elements of $\Sigma$).
@@
In this subsection, we introduce partial actions on sets, discuss the concept of globalization, and provide illustrative examples.
We begin by defining a partial action. From now on, $X$ will denote a set.
@@
Consequently, ($(\{_{s}X\}_{s\in S},\{\alpha_{s}\}_{s\in S})$ is a global action if and only if it satisfies (G1) and (ii).
@@
${{}_{s}X}\cap X_{t}=\alpha_{t}({{}_{t}X}\cap{{}_{st}X})=\alpha_{t}({{}_{t}X})=X_{t}\quad\therefore\quad X_{t}\subseteq{{}_{s}X}.$
Now, (P2) is $(\alpha_{s}\circ\alpha_{t})(x)=\alpha_{st}(x)$, for every $x\in{{}_{st}X}={{}_{t}X}$, so $\alpha_{s}\circ\alpha_{t}=\alpha_{st}$ as functions. Conversely, assume (ii). By hypothesis, (G2) holds, and the condition $\alpha_{s}\circ\alpha_{t}=\alpha_{st}$ implies (P2). Furthermore, we have $\alpha_{t}^{-1}({{}_{s}X}\cap X_{t})=\alpha_{t}^{-1}(X_{t})={{}_{t}X}={tX}\cap{{st}X}$, so (P1) holds. ∎
@@
Let $(\alpha,X)$ be a partial action satisfying (G1). If $S^{s}\neq\emptyset$, for every $s\in S$, then $(\alpha,X)$ satisfies (G2). Therefore $(\alpha,X)$ is a global action of $S$.
@@
Example 2.12 illustrates that a semigroupoid always acts globally on itself via left multiplication. When $S$ is specifically a semigroup or a category, for every $t\in S$, there exists some $s\in S$ such that $(s,t)\in S^{(2)}$. In the case of a category, we can choose $s=R(t)$, the left identity of $t$, and for a semigroup, we can choose any $s\in S$. Consequently, $S=\bigcup_{s\in S}S^{s}=\bigcup_{s\in S}{{}_{s}X}$, and the left multiplication action is non-degenerate. This is not always the case for a semigroupoid. For instance, the left multiplication of the semigroupoid $S=\{s,t,st\}$, presented in Example 2.6, satisfies $\bigcup_{s\in S}({{}_{s}X}\cup X_{s})=\{t,st\}\neq S$.
@@
${{}_{s_{1}}X}$ =$=\{1,2\},$ $\alpha_{s_{1}}(1,2)$ $=(2,3),$ ${{}_{s_{2}}X}$ =$=\{1\},$ $\alpha_{s_{2}}(1)$ $=(2),$
${{}_{t_{1}}X}$ =$=\{2\},$ $\alpha_{t_{1}}(2)$ $=(1),$ ${{}_{t_{2}}X}$ =$=\{1,3\},$ $\alpha_{t_{2}}(1,3)$ $=(1,3),$
${{}_{0}X}$ =$=\{1,2\},$ $\alpha_{0}(1,2)$ $=(2,2).$
@@
In this section, we present the construction of partial actions by restricting global ones and motivate our definition of universal globalization.
@@
=$\overset{(\ast\ast)}{=}\beta_{t}^{-1}{\Big{(}}\beta_{s}^{-1}(X\cap Y_{s})\cap X\cap Y_{t}{\Big{)}}\cap X$
@@
=$\overset{(\ast)}{=}\alpha_{t}^{-1}({{}_{s}X}\cap X_{t}).$
So condition (P1) holds. Since $\alpha_{s}=\beta_{s}|_{{}_{s}X}$, condition (P2) is easily verified.
[…]
@@
In many contexts, a partial action $(\alpha,X)$ is called a $S$-act, and a morphism of partial actions $(\alpha,X)\to(\beta,Y)$ is called a $S$-function. This is the convention, for instance, in [17], [21] and [8]. Here we chose to adopt the terminology presented by Abadie in [1] and his subsequential works.
@@
(i) Suppose that $\varphi\colon(\alpha,X)\to(\beta,Y)$ is an isomorphism of $\mathcal{A}_{p}(S)$. Then there exists a morphism of partial actions $\psi\colon(\beta,Y)\to(\alpha,X)$ such that $\varphi\circ\psi=id_{(\alpha,X)}$ and $\psi\circ\varphi=id_{(\beta,Y)}$. In particular, we have $\psi=\varphi^{-1}$ as a function, so $\varphi$ is bijective. Now, given $s\in S$, we calculate
@@
For the reverse inclusion, let $x\in{{}_{s}X}$. Then $\varphi(x)\in{{}_{s}Y}$ and $\beta_{s}(\varphi(x))=\varphi(\alpha_{s}(x))\in\varphi(X)$. So every isomorphism in $\mathcal{A}_{p}(S)$ is a bijective embedding.
@@
So $\varphi\colon(\alpha,X)\to(\beta,Y)$ is a morphism of partial action. Furthermore, we have
@@
Let $(\alpha,X)$ be a partial action of $S$. A globalization of $(\alpha,X)$ is a triple $(\delta,\beta,Y)$ consisting of a global action $(\beta,Y)$ and an embedding $\delta\colon X\to Y$.
@@
Then each pair $(\alpha^{n},X^{n})$ is a global action of $\mathcal{S}$, and each inclusion $i^{n}\colon X^{1}\to X^{n}$ is an embedding. Hence, each triple $(\alpha^{n},X^{n},i^{n})$ is a globalization of $(\alpha^{1},X^{1})$. However, there is no bijection $X^{i}\to X^{j}$ when $i\neq j$. Therefore $(\alpha^{1},X^{1})$ has infinitely many non-isomorphic globalizations.
Our interest is to define a class of globalizations in a way that, for each partial action, there is at most one globalization (up to isomorphism) in that class. However, there are multiple ways to define such class. For instance, in the context of semigroups, [17, Theorem 3.5] and [17, Theorem 4.7] present non-isomorphic globalizations that are unique in their respective class. Here we consider the class of universal globalizations, which appear in [1, Definition 2.24] as enveloping actions.
@@
Let $(\alpha,X)$ be a partial action. A universal globalization of $(\alpha,X)$ is a globalization $(\delta,\beta,Y)$ of $(\alpha,X)$ such that, for every morphism $\varphi\colon(\alpha,X)\to(\theta,Z)$, where $(\theta,Z)$ is a global action, there is an unique morphism $\Phi\colon(\beta,Y)\to(\theta,Z)$ such that $\Phi\circ\delta=\varphi$. That is, the following diagram commutes:
$(\alpha,X)$$(\beta,Y)$$(\theta,Z)$$\delta$$\varphi$$\Phi$
@@
Let $(\alpha,X)$ be a partial action of $S$ and suppose that $(\beta,Y,\delta)$ and $(\theta,Z,\gamma)$ are universal globalizations of $(\alpha,X)$. Then there is an isomorphism of partial actions $(\beta,Y)\to(\theta,Z)$.
@@
Since $(\beta,Y,\delta)$ is a universal globalization of $(\alpha,X)$, the function $\delta\colon X\to Y$ is, in particular, a morphism of partial actions. As $(\beta,Y)$ is a global action and $(\theta,Z,\gamma)$ is a universal globalization, there must be a morphism of partial actions $f\colon(\theta,Z)\to(\beta,Y)$ such that $\delta=f\circ\gamma$.
Interchanging the roles of $(\beta,Y,\delta)$ and $(\theta,Z,\gamma)$ above, we obtain a morphism of partial action $g\colon(\beta,Y)\to(\theta,Z)$ such that $\gamma=g\circ\delta$. Now, notice that $f\circ g\colon(\beta,Y)\to(\beta,Y)$ satisfies
$(f\circ g)\circ\delta=f\circ\gamma=\delta=id_{Y}\circ\delta.$
Since $(\beta,Y,\delta)$ is a universal globalization, the morphism $F\colon(\beta,Y)\to(\beta,Y)$ such that $F\circ\delta=\delta$ must be unique. Therefore $f\circ g=id_{Y}$. Analogously, the composition $g\circ f\colon(\theta,Z)\to(\theta,Z)$ satisfies
$(g\circ f)\circ\gamma=g\circ\delta=\gamma=id_{Z}\circ\gamma.$
Since $(\theta,Z,\gamma)$ is a universal globalization, it must be $g\circ f=id_{Z}$. Therefore $f=g^{-1}$. By definition, both $f$ and $g$ are morphisms of partial actions. Thus, $g\colon(\beta,Y)\to(\theta,Z)$ is an isomorphism. ∎
It is easy to see that, if $(\alpha,X)$ is a global action, then $(\alpha,X,id_{X})$ is the universal globalization of $(\alpha,X)$. In this case, given a morphism $\varphi\colon(\alpha,X)\to(\beta,Y)$, we can chose $\Phi=\varphi$. Therefore, in Example 2.20, $(\alpha^{1},X^{1},i^{1})$ is the unique universal globalization of $(\alpha^{1},X^{1})$.
Before we present the construction of a universal globalization, we want to point it out that, if universal globalizations do exist, then they can be characterized as the initial objects of a suitable category. For it, let $\mathcal{C}$ and $\mathcal{D}$ be categories. Recall that a functor $F\colon\mathcal{C}\to\mathcal{D}$ is a pair of functions $F_{0}\colon\mathcal{C}_{0}\to\mathcal{D}_{0}$ and $F_{1}\colon\mathcal{C}_{1}\to\mathcal{D}_{1}$ such that:

### v2

@@ Abstract.
We introduce partial semigroupoid actions on sets and demonstrate that each such action admits a universal globalization. Our construction extends the universal globalization for partial category actions given by P. Nystedt (Lundström) and the tensor product globalization for strong partial semigroup actions given by G. Kudryavtseva and V. Laan, thereby unifying the theory of partial actions for both categories and semigroups.
@@ 1. Introduction
Partial group actions on sets were introduced by R. Exel in [11], where he characterizes a class of $C^{\ast}$-algebras as crossed products by partial actions. Examples of partial group action on sets can be easily obtained from a restriction of a global group action to a subset of a given set. In [1, Teorema 4.23], F. Abadie proved that any partial group action on a set can indeed be viewed as a restriction of a unique up to isomorphism global one. This global action was later called universal globalization, due to its universal property. In [17, Theorem 3.4], J. Kellendonk and M. Lawson proved that universal globalizations are initial objects in a certain category. Partial actions and the globalization problem have been investigated in many other cases, such as partial actions of monoids ([16], [18] and [23]), semigroups [19], restriction semigroups [15], ordered groupoids ([14] and [21]), groupoids [3], inverse categories [2], categories [24], and inverse semigroupoids [8]. The reader can also find more information on partial actions in the surveys [5], [9] and [10], and in Exel’s book [13].
@@
As an algebraic structure, a semigroupoid is a set equipped with a partially defined associative operation. It is worth mentioning that there is another (more restrictive) notion of semigroupoid due to B. Tilson [25]. The latter can be seen as a directed graph equipped with a partially defined associative operation, and therefore it contains the classes of categories and semigroups. However, the class of Exel’s semigroupoids is strictly larger than the class of Tilson’s semigroupoids (also referred to as semicategories or categorical semigroupoids). Several properties of semigroupoids and graphed semigroupoids can be found in the work by L. Cordeiro in [6].
In this paper, we are interested in studying the globalization problem for partial semigroupoid actions on sets. Our method generalizes the work of P. Nystedt (Lundström) in [24], whose construction of a universal globalization for partial category actions traces back to F. Abadie’s approach in [1, Theorem 1.1] and Kellendonk and Lawson’s approach in [17, Theorem 3.4]. We also compare our results with the tensor product globalization for partial semigroup actions constructed by G. Kudryavtseva and V. Laan in [19], concluding that our universal globalization coincides with the tensor product globalization when the partial semigroup action is non-degenerate.
Partial category and semigroup actions extend the notion of partial monoid actions studied by M. Megrelishvili and L. Schröder [23], which in turn generalize the classical partial group actions of Exel [11] and of Kellendonk and Lawson [17]. An alternative line of generalization, due to P. Demeneghi and F. Tasca [8], addresses the globalization problem for partial actions of inverse semigroupoids. In the final section of this work, we discuss in detail the differences between the semigroupoid and the inverse semigroupoid settings.
The paper is organized as follows. Section 2 recalls the definition of semigroupoids and introduces the notion of partial actions, together with the construction of the standard restriction of a global action to a subset and the definition of universal globalization. Section 3 develops the construction of a universal globalization for any given partial action. In Section 4, we discuss the relationship between our construction and those previously established for semigroups and categories. Finally, Section 5 highlights the differences between partial actions of semigroupoids and those of inverse semigroupoids.
@@ 2. Global and Partial Actions
For our purposes, by a semigroupoid we mean an Exel’s semigroupoid, which is a more general notion than that of Tilson’s semigroupoid.
@@
In other words, $S^{s}$ is the set of elements in $S$ that are right composable with $s$, while $^{s}S$ is the set of elements in $S$ that are left composable with $s$. Due to the associativity of $S$, we have $S^{t}=S^{st}$ and ${{}^{s}S}={{}^{st}S}$, whenever $(s,t)\in S^{(2)}$. A semigroupoid is called categorical if, for every $s,t\in S$, the sets $S^{s}$ and $S^{t}$ are either disjoint or equal. By [6, Proposition 2.14], a semigroupoid is categorical if and only if, for all $s,t\in S$, the sets $^{s}S$} and $^{t}S$} are either disjoint or equal.
The class of semigroupoids generalizes two relevant algebraic structures: the semigroups, which can be seen as the semigroupoids whose composition is totally defined; and categories, which can be seen as semigroupoids with local identities. We present the definition of those classes for future reference.
@@
By definition, semigroups are precisely semigroupoids such that $S^{(2)}=S\times S$. Furthermore, every semigroup is a categorical semigroupoid, since $\mathcal{S}^{s}=\mathcal{S}^{t}$ for every pair $s,t\in\mathcal{S}$.
@@
Notice that, if $1_{e}\circ 1_{f}$ is defined, then $1_{e}=1_{f}$ and, therefore, $e=D(1_{e})=D(1_{f})=f$. Consequently, $e\mapsto 1_{e}$ is an injection and, hence, we can assume $\mathcal{C}_{0}\subseteq\mathcal{C}_{1}=:\mathcal{C}$. Furthermore, for every $f\in\mathcal{C}_{1}$, the composition $1_{R(f)}\circ f$ is defined. Hence, if $f$ is an identity, then $f=1_{R(f)}\in\mathcal{C}_{0}$, which shows that $\mathcal{C}_{0}$ is precisely the set of identities of $\mathcal{C}_{1}$.
A category can be seen as a semigroupoid such that $\mathcal{C}^{(2)}=\{(f,g)\colon D(f)=R(g)\}\subseteq\mathcal{C}\times\mathcal{C}$, and such that every element $f\in\mathcal{C}$ has a left identity $R(f)$ and a right identity $D(f)$. Moreover, every category is a categorical semigroupoid. In fact, for every $f\in\mathcal{C}$ we have:
@@
[25, Appendix B] A graphed semigroupoid (also called Tilson semigroupoid) is a quintuple $(G_{0},G_{1},d,r,\star)$, where:
@@
Markov semigroupoids constitute a class of semigroupoids that contains the path categories of directed graphs and some semigroupoids that cannot be realized as graphed semigroupoids. They are constructed in the following way. Let $\Sigma$ be a set and $A=(A(x,y))_{x,y\in\Sigma}$ be a 0-1 matrix. Assume that:
@@
Let $\Sigma=\{s,t\}$ and $A$ be the $2\times 2$ matrix such that $A(s,t)=1$ and $A(x,y)=0$, for all $(x,y)\neq(s,t)$. The Markov semigroupoid associated to $(\Sigma,A)$ is the set $S=\{s,t,st\}$, where the only composition defined is $s\star t=st$. This Markov semigroupoid can be endowed with a graph structure $(S_{0},S_{1},d,r)$, in the following way: $S_{0}$ consists of three vertex, say $S_{0}=\{e_{1},e_{2},e_{3}\}$, $S_{1}=S$, and the functions $d,r\colon S\to S_{0}$ are given by $d(t)=d(st)=e_{1}$, $d(s)=r(t)=e_{2}$ and $r(s)=r(st)=e_{3}$. The semigroupoid $S$ associated with this graph structure is shown in the following diagram.
$\overset{e_{1}}{\bullet}$$\overset{e_{2}}{\bullet}$$\overset{e_{3}}{\bullet}$ttssstst
@@
We now show that $S$ is not a Markov semigroupoid either. Indeed, if it were Markov, then $\Sigma$ would be the set of sequences of length one in $S$, that is, the set of elements that cannot be written as as compositions of other elements. In this case, $\Sigma=\{s_{1},s_{2},t_{1},t_{2}\}$. However, we would then have $s_{1}\star t_{1}=(s_{1},t_{1})\neq(s_{1},t_{2})=s_{1}\star t_{2}$, which contradicts the definition of the product $s_{i}\star t_{j}=0$ for all $(s_{i},t_{j})\in S^{(2)}$. Hence, $S$ is not a Markov semigroupoid.
@@
In this subsection, we introduce the notion of partial actions of semigroupoids on sets, discuss the concept of globalization, and provide illustrative examples.
We begin by defining what a partial action is. From this point on, $X$ denotes a set. Given a set $Y$, a function $f\colon Y\to X$, and a subset $A\subseteq X$, we denote by $f^{-1}(A)$ the inverse image of $A$ by $f$. That is, $f^{-1}(A)=\{x\in Y\colon f(x)\in A\}$.
@@
Consequently, $(\{_{s}X\}_{s\in S},\{\alpha_{s}\}_{s\in S})$ is a global action if and only if it satisfies (ii), for all $(s,t)\in S^{(2)}$, and (G1).
@@
${{}_{s}X}\cap X_{t}=\alpha_{t}({{}_{t}X}\cap{{}_{st}X})=\alpha_{t}({{}_{t}X})=X_{t}.$
Therefore, $X_{t}\subseteq{{}_{s}X}$. Now, (P2) is $(\alpha_{s}\circ\alpha_{t})(x)=\alpha_{st}(x)$, for every $x\in{{}_{st}X}={{}_{t}X}$, thus $\alpha_{s}\circ\alpha_{t}=\alpha_{st}$ as functions. Conversely, assume (ii). By hypothesis, (G2) holds, and the condition $\alpha_{s}\circ\alpha_{t}=\alpha_{st}$ implies (P2). Furthermore, we have $\alpha_{t}^{-1}({{}_{s}X}\cap X_{t})=\alpha_{t}^{-1}(X_{t})={{}_{t}X}={{}_{t}X}\cap{{}_{st}X}$. This proves (P1).
In conclusion, if (ii) is satisfied for all $(s,t)\in S^{(2)}$, then the pair $(\{_{s}X\}_{s\in S},\{\alpha_{s}\}_{s\in S})$ satisfies (P1), (P2) and (G2). If it also satisfies (G1), then it is a global action by definition. ∎
@@
Let $\alpha$ be a partial action satisfying (G1). If $S^{s}\neq\emptyset$, for every $s\in S$, then $(\alpha,X)$ satisfies (G2). Therefore $\alpha$ is a global action of $S$.
@@
Example 2.12 illustrates that a semigroupoid always acts globally on itself via left multiplication. When $S$ is specifically a semigroup or a category, for every $t\in S$, there exists some $s\in S$ such that $(s,t)\in S^{(2)}$. In the case of a category, we can choose $s=R(t)$, the left identity of $t$, and for a semigroup, we can choose any $s\in S$. Consequently, $S=\bigcup_{s\in S}S^{s}=\bigcup_{s\in S}{{}_{s}X}$, and the left multiplication action is non-degenerate. This is not always the case for a semigroupoid. For instance, the left multiplication of the semigroupoid $S=\{s,t,st\}$, as defined in Example 2.7, satisfies $\bigcup_{s\in S}({{}_{s}X}\cup X_{s})=\{t,st\}\neq S$.
@@
${{}_{s_{1}}X}$ $=\{1,2\},$ $\alpha_{s_{1}}(1)$ $=2,\ \alpha_{s_{1}}(2)=3,$ ${{}_{s_{2}}X}$ $=\{1\},$ $\alpha_{s_{2}}(1)$ $=2,$
${{}_{t_{2}}X}$ $=\{1,3\},$ $\alpha_{t_{2}}(1)$ $=1,\ \alpha_{t_{2}}(3)=3,$ ${{}_{t_{1}}X}$ $=\{2\},$ $\alpha_{t_{1}}(2)$ $=1,$
${{}_{0}X}$ $=\{1,2\},$ $\alpha_{0}(1)$ $=2,\ \alpha_{0}(2)=2.$
@@
In this section, we present the construction of partial actions by restricting global ones, and motivate our definition of universal globalization.
@@
$\overset{(\ast\ast)}{=}\beta_{t}^{-1}{\Big(}\beta_{s}^{-1}(X\cap Y_{s})\cap X\cap Y_{t}{\Big)}\cap X$
@@
$\overset{(\ast)}{=}\alpha_{t}^{-1}({{}_{s}X}\cap X_{t}).$
This proves that condition (P1) holds. Since $\alpha_{s}=\beta_{s}|_{{}_{s}X}$, condition (P2) is easily verified.
It remains to prove ($\ast$) and ($\ast\ast$). For the first one, if $x\in\beta_{t}^{-1}({{}_{s}X}\cap Y_{t})\cap X\subseteq X\cap{{}_{t}Y}$, then $\beta_{t}(x)\in{{}_{s}X}\cap Y_{t}\subseteq X$. Since $x\in{{}_{t}X}$, we have $\beta_{t}(x)=\alpha_{t}(x)\in X_{t}$. Therefore,
@@
In many contexts, a partial action $(\alpha,X)$ is called a $S$-act, and a morphism of partial actions $(\alpha,X)\to(\beta,Y)$ is called a $S$-function. This is the convention, for instance, in [19], [24] and [8]. Here we chose to adopt the terminology of Abadie from [1] and his subsequential works.
@@
(i) Suppose that $\varphi\colon(\alpha,X)\to(\beta,Y)$ is an isomorphism of $\mathcal{A}_{p}(S)$. Then there exists a morphism of partial actions $\psi\colon(\beta,Y)\to(\alpha,X)$ such that $\varphi\circ\psi=id_{(\alpha,X)}$ and $\psi\circ\varphi=id_{(\beta,Y)}$. In particular, we have $\psi=\varphi^{-1}$ as a function, thus $\varphi$ is bijective. Now, given $s\in S$, we calculate
@@
For the reverse inclusion, let $x\in{{}_{s}X}$. Then $\varphi(x)\in{{}_{s}Y}$ and $\beta_{s}(\varphi(x))=\varphi(\alpha_{s}(x))\in\varphi(X)$. Thus, every isomorphism in $\mathcal{A}_{p}(S)$ is a bijective embedding.
@@
Thus, $\varphi\colon(\alpha,X)\to(\beta,Y)$ is a morphism of partial action. Furthermore, we have
@@
Let $(\alpha,X)$ be a partial action of $S$. A globalization of $(\alpha,X)$ is a triple $(\iota,\beta,Y)$ consisting of a global action $(\beta,Y)$ and an embedding $\iota\colon X\to Y$.
@@
Then each pair $(\alpha^{n},X^{n})$ is a global action of $\mathcal{S}$, and each inclusion $i^{n}\colon X^{1}\to X^{n}$ is an embedding. Hence, each triple $(i^{n},\alpha^{n},X^{n})$ is a globalization of $(\alpha^{1},X^{1})$. However, there is no bijection $X^{i}\to X^{j}$ when $i\neq j$. Therefore $(\alpha^{1},X^{1})$ has infinitely many non-isomorphic globalizations.
Our interest is to define a class of globalizations in a way that, for each partial action, there exists at most one globalization (up to isomorphism) in that class. However, there are multiple ways to define such class. For instance, in the context of semigroups, [19, Theorem 3.5] and [19, Theorem 4.7] present non-isomorphic globalizations that are unique in their respective class. Here we consider the class of universal globalizations, which appear in [1, Definition 2.24] as enveloping actions.
@@
Let $(\alpha,X)$ be a partial action. A universal globalization of $(\alpha,X)$ is a globalization $(\iota,\beta,Y)$ of $(\alpha,X)$ such that, for every morphism $\varphi\colon(\alpha,X)\to(\theta,Z)$, where $(\theta,Z)$ is a global action, there exists a unique morphism $\Phi\colon(\beta,Y)\to(\theta,Z)$ such that $\Phi\circ\iota=\varphi$. That is, the following diagram commutes:
$(\alpha,X)$$(\beta,Y)$$(\theta,Z)$$\iota$$\varphi$$\Phi$
@@
Let $(\alpha,X)$ be a partial action of $S$ and suppose that $(\iota,\beta,Y)$ and $(\gamma,\theta,Z)$ are universal globalizations of $(\alpha,X)$. Then there is an isomorphism of partial actions $(\beta,Y)\to(\theta,Z)$.
@@
Since $(\iota,\beta,Y)$ is a universal globalization of $(\alpha,X)$, the function $\iota\colon X\to Y$ is, in particular, a morphism of partial actions. As $(\beta,Y)$ is a global action and $(\gamma,\theta,Z)$ is a universal globalization, there must be a morphism of partial actions $f\colon(\theta,Z)\to(\beta,Y)$ such that $\iota=f\circ\gamma$.
Interchanging the roles of $(\iota,\beta,Y)$ and $(\gamma,\theta,Z)$ above, we obtain a morphism of partial action $g\colon(\beta,Y)\to(\theta,Z)$ such that $\gamma=g\circ\iota$. Now, notice that $f\circ g\colon(\beta,Y)\to(\beta,Y)$ satisfies
$(f\circ g)\circ\iota=f\circ\gamma=\iota=id_{Y}\circ\iota.$
Since $(\iota,\beta,Y)$ is a universal globalization, the morphism $F\colon(\beta,Y)\to(\beta,Y)$ such that $F\circ\iota=\iota$ must be unique. Therefore $f\circ g=id_{Y}$. Analogously, the composition $g\circ f\colon(\theta,Z)\to(\theta,Z)$ satisfies
$(g\circ f)\circ\gamma=g\circ\iota=\gamma=id_{Z}\circ\gamma.$
Since $(\gamma,\theta,Z)$ is a universal globalization, it must be $g\circ f=id_{Z}$. Therefore $f=g^{-1}$. By definition, both $f$ and $g$ are morphisms of partial actions. Thus, $g\colon(\beta,Y)\to(\theta,Z)$ is an isomorphism. ∎
It is easy to see that, if $(\alpha,X)$ is a global action, then $(id_{X},\alpha,X)$ is the universal globalization of $(\alpha,X)$. In this case, given a morphism $\varphi\colon(\alpha,X)\to(\beta,Y)$, we can chose $\Phi=\varphi$. Therefore, in Example 2.20, $(i^{1},\alpha^{1},X^{1})$ is the unique universal globalization of $(\alpha^{1},X^{1})$.
Before constructing a universal globalization, we want to point out that, if universal globalizations exist, then they can be characterized as the initial objects of a suitable category. To prove this, let $\mathcal{C}$ and $\mathcal{D}$ be categories. Recall that a covariant functor $F\colon\mathcal{C}\to\mathcal{D}$ is a pair of functions $F_{0}\colon\mathcal{C}_{0}\to\mathcal{D}_{0}$ and $F_{1}\colon\mathcal{C}_{1}\to\mathcal{D}_{1}$ such that: